You are given the following information on two salespeople. Salesperson A earns a $22,000 salary; salesperson B earns $23,000. Both earn 1 percent commission. Advertising costs $3 per unit. Shipping expenses are $2 per unit. Order processing costs total $1 per order. Travel expenses amount to $0.50 per call. Your task is to calculate the contribution to profit (loss) made by each salesperson.

Salesperson

Number of Calls

Number of orders

Units sold

Total Sales

Cost of Goods Sold

A

200

250

15,000

$750,000

$600,000

B

295

230

18,000

900,000

720,000
Answer #1

Salesman 1

Salary=$22,000

1% commission for $750,000 sales= 750,000*(1/100)=$7500

Advertisement expense=Adv cost*number of units= 3*15000= $45,000

Total shipping expenses= Shipping expense per unit* number of units sold= 2*15000=$30,000

Total order processing cost= order processing cost per unit* number of units sold= 1*15000=$15,000

Total call expenses= travel expense*number of calls made= 0.5*200= $100

Total expense= 22,000+7,500+45,000+30,000+15,000+100= $119,600

Contribution to profit = Total sales amount-(cost of goods sold+ total expenses)= 900,000-(720,000+140,147.5) = $30,400

Salesman 2

Salary= $23000

Commission= 1% of 900,000= $9,000

Advertisement expense=Adv cost*number of units= 3*18000= $54,000

Total shipping expenses= Shipping expense per unit* number of units sold= 2*18000=$36,000

Total order processing cost= order processing cost per unit* number of units sold= 1*18000=$18,000

Total call expenses= travel expense*number of calls made= 0.5*295= $147.5

Total expense= 23,000+9,000+54,000+36,000+18,000+147.5= $140,147.5

Contribution to profit = Total sales amount-(cost of goods sold+ total expenses)=900,000-(720,000+140147.5)= $39,852.5